The SP Corporation makes 32,000 motors to be used in the production of its sewing machines. The average cost per motor at this level of activity is:

Direct materials $ 9.10
Direct labor $ 8.10
Variable manufacturing overhead $ 3.25
Fixed manufacturing overhead $ 4.20

An outside supplier recently began producing a comparable motor that could be used in the sewing machine. The price offered to SP Corporation for this motor is $22.75. If SP Corporation decides not to make the motors, there would be no other use for the production facilities and none of the fixed manufacturing overhead cost could be avoided. Direct labor is a variable cost in this company. The annual financial advantage (disadvantage) for the company as a result of making the motors rather than buying them from the outside supplier would be:

Multiple Choice

  • ($60,800)

  • $177,600

  • 73,600

  • $134,400
Answer #1

Financial advantage (Disadvantage) of making rather than buying

Cost of making = (9.1+8.1+3.25)*32,000 = 654,400

Cost of buying = 22.75*32,000 = 728,000

= 728,000 - 654,400

= 73,600

Option C